# Neutral-current background induced by atmospheric neutrinos at large liquid-scintillator detectors: III. Comprehensive prediction for low energy neutrinos

### v2

@@ 2.3 GEANT4-based Detector Simulation
GEANT4 (version 4.10.p02) is utilized to simulate the propagation of final-state particles resulting from -12C NC interactions within the LS detector. The simulation employs the QGSPBERT physics list, with the Quark Gluon String (QGS) model applied for hadronic interactions at higher energies. For hadronic interactions at the lowest energy range (below 70 MeV) and nuclear deexcitation coupled with the higher-energy QGS model, GEANT4’s precompound (P) model is employed. Additionally, the Bertini Cascade (BERT) model is applied for hadronic interactions at the lower energy range. The High-Precision neutron (HP) model is used for neutron elastic and inelastic interactions below 20 MeV, ensuring high accuracy in modeling these interactions. The detector simulation incorporates the decay processes of unstable residual nuclei, utilizing information such as decay types, endpoints, and lifetimes sourced from the nuclear database [56].
@@ 3.2.1 Production rates of final-state particles
Initially, we calculate the production rates of relevant final-state particles from both QE and . However, observation reveals that GENIE (3.0.6) generates nucleons with zero kinetic energy, a phenomenon also identified in the older GENIE model [45]. Our investigation extends to the newer version of GENIE (3.4.0), where the issue of final-state nucleons with zero kinetic energies persists. Consultation with GENIE’s authors clarifies that this problem stems from the handling of binding energy during the NCQE process in GENIE. Note that the most recent GENIE versions are anticipated to improve the handling of binding energy in the modeling of the FSI processes. To address the issue, their recommendation involves marking a nucleon as intermediate if it lacks sufficient kinetic energy to leave a nucleus (e.g., its kinetic energy inside the nucleus is lower than the binding energy), with all its energy-momentum absorbed by the hadronic blob. Following these suggestions, the final-state information from the new GENIE models (i.e., Model-G for 2, 3, 4) is modified. To distinguish models subjected to this operation from those that are not, the modified models are re-labeled as Model-G for 2’, 3’, 4’.
@@
“New GENIE’ model variations”: the model variations among Model-G for 2’, 3’, 4’, which are used to demonstrate the impact of modifications in the new GENIE models, particularly addressing the issue of zero kinetic energy nucleons production in final-state results.
@@
When excluding nucleons with zero kinetic energies in the final-state, Model-G2’/4’ (N) exhibit a 25% lower rate of neutrons compared to the associated models. Similarly, Model-G3’ (A) shows a 10% lower rate of neutrons in comparison. Therefore, in the new GENIE’ models, Model-G2’ still exhibits the lowest rates, while Model-G3’ shows the highest rates.
@@
Fig. 5 presents the production rates of primary final-state particles, such as and , for kinetic energy. The dominant contributors to production are the QE and . To enhance clarity, nucleons with zero kinetic energy are excluded in Fig. 5. This operation results in kinetic energy spectra in Model-G 2, 3, 4 aligned with those from the corresponding Model-G 2’, 3’, 4’ in the left panels of Fig. 5, illustrating the spectra of and from NC interactions without deexcitation. When focusing on nucleon productions excluding those from deexcitation, we observe the same conclusions as depicted in Fig. 4. Furthermore, the impact of different nuclear models on production rates falls within the (0 20) MeV range in kinetic energy, as evident in the comparison between Model-G2 and Model-G4, or Model-N6 and Model-N7. Similarly, the effects of different FSI models on production rates are within the (0 40) MeV range in kinetic energy, observed in comparisons such as Model-G2 versus Model-G3, or Model-G2 versus Model-N6. By comparing the left and right panels of Fig. 5, it becomes apparent that the enhanced production rates of and from deexcitation occur within the (0 20) MeV range of kinetic energy. As expected, due to slight differences in the excited residual nucleus, the kinetic energy spectra of Model-G 2’, 3’, 4’ exhibit minor variations from the corresponding Model-G 2, 3, 4.
@@ 3.2.2 Neutron multiplicity
In Fig. 6, we examine the neutron multiplicity in the final-state of NC interactions involving -12C. Considering that the predominant neutron multiplicities are below three, the event rates for neutron multiplicities less than three are highlighted in the subpanels of the upper panels of Fig. 6. Due to the significant similarity in distributions between Model-G2 and Model-G4, as well as between Model-G2’ and Model-G4’, Model-G4/4’ is excluded from the illustrations. Excluding neutrons from the deexcitation process, the events with neutron multiplicities less than three account for approximately 91% in the new GENIE models using the N FSI model (i.e., Model-G2/4) and A FSI model (i.e., Model-G3). This calculation includes neutrons with zero kinetic energies in the final-state. If we exclude these neutrons in Model-G 2’, 3’, 4’, the events with neutron multiplicities less than three increase to 95% in Model-G2’/4’, while the fraction remains unchanged in Model-G3’. This difference arises because neutrons with zero kinetic energies from the N FSI model impact all neutron multiplicities, whereas those from the A FSI model only affect neutron multiplicities less than three. Compared to GENIE models, NuWro models produce around 99% events with neutron multiplicities less than three. For comparison purposes, the middle panels of Fig. 6 display the ratio of results related to multiplicities less than three to those in Model-G2. The neutrons induced by the deexcitation process increase the multiplicity. The parameter, , defined as the ratio of event rates for NC events with deexcitation to those for NC events without deexcitation, concerning multiplicities less than three is illustrated in the lowest panel of Fig. 6.
@@
Upon scrutinizing model variations, the maximum difference is observed in NC events associated with 10B for the new GENIE model variations. However, considering that nucleons with zero kinetic energy remain within the nuclei, the disparity in NC events with 10B is significantly mitigated in new GENIE’ model variations, resulting in lower event rates. Meanwhile, the event rate for NC events with 11C sees a noteworthy increase in new GENIE’ models. Regarding new NuWro model variations, the primary distinction arises in NC events with 11C, while other channels exhibit similarity in the absence of deexcitation scenarios. Notably, Model-N6 with the LFG nuclear model produces substantially higher event rates for NC events with 11C.
@@ 4 NC Background Predictions at Low Visible Energies
Fig. 10 presents event rates with different final-state nuclei across distinct prompt energy ranges of (0.7 100) MeV, (0.7 12) MeV, and (12 30) MeV, where the last two energy regions are associated with the detection regions of reactor neutrino and DSNB signals, respectively. The predominance of the NC background with 11C is evident. The deexcitation effect minimally impacts the overall event rate, causing a reduction in rates for dominant channels with 11C and 10B, while increasing rates for channels with lighter nuclei such as 9Be, 6Li and 8Be. When comparing the new and old neutrino interaction models, the rates of the new models do not surpass those of the old models, with only slightly higher rates observed for new GENIE’ models in the (0.7 12) MeV energy range.
Tab. 2 summarizes the average rates of the NC background rates, with the model uncertainty represented by the 1 deviation, for both old and new neutrino interaction models. In this table, the old models considered are Model-G1 and Model-N, where 1, 2, 3, 4, 5. Due to a strong correlation between the new GENIE and GENIE’ models, we opt to use the new GENIE’ model calculations instead of the new GENIE models for the prediction. Consequently, the results from the new models in Tab. 2 include Model-G, 2’, 3’, 4’, and Model-N, 6, 7. To investigate the impact of the deexcitation on the final NC background rate, both results without and with deexcitation are presented in Tab. 2. It’s important to note that while the deexcitation model used in this work still requires constraints from experimental data, the NC background rate with deexcitation is considered as the nominal result, as deexcitation always occurs in a realistic condition. In the prompt range of (0.7 100) MeV, the average NC background rates of the new models are lower than those of the old models for both scenarios, with and without the deexcitation process. This difference arises from the fact that the new models no longer incorporate the model with 1.35 GeV, which previously led to a higher NC background rate in the higher energy range. The deexcitation effect results in an increase in the average NC background rates by around a factor of 1.06 and 1.09 for old models and new models in this energy range, respectively. When considering the deexcitation, the associated uncertainty arising from model variation is around 16% and 14% for old and new models, respectively. In the (0.7 12) MeV and (12 30) MeV energy ranges, we will discuss the results related to the detection of reactor and DSNB in the LS detectors.
@@ Acknowledgements
The authors would like to thank Wan-Lei Guo and Xian-Guo Lu for helpful discussions. We also thank Davide Basilico and Liang Zhan for carefully reading the manuscript and useful comments. This work is supported in part by National Natural Science Foundation of China under Grant Nos. 12075255, 11835013, 12125506.

### v3

@@ 2.3 GEANT4-based Detector Simulation
GEANT4 (version 4.10.p02) is utilized to simulate the propagation of final-state particles resulting from -12C NC interactions within the LS detector. The simulation employs the QGSPBERT physics list, with the Quark Gluon String (QGS) model applied for hadronic interactions at higher energies. For hadronic interactions at the lowest energy range (below 70 MeV) and nuclear deexcitation coupled with the higher-energy QGS model, GEANT4′s precompound (P) model is employed. Additionally, the Bertini Cascade (BERT) model is applied for hadronic interactions at the lower energy range. The High-Precision neutron (HP) model is used for neutron elastic and inelastic interactions below 20 MeV, ensuring high accuracy in modeling these interactions. The detector simulation incorporates the decay processes of unstable residual nuclei, utilizing information such as decay types, endpoints, and lifetimes sourced from the nuclear database [56].
@@ 3.2.1 Production rates of final-state particles
Initially, we calculate the production rates of relevant final-state particles from both QE and . However, observation reveals that GENIE (3.0.6) generates nucleons with zero kinetic energy, a phenomenon also identified in the older GENIE model [45]. Our investigation extends to the newer version of GENIE (3.4.0), where the issue of final-state nucleons with zero kinetic energies persists. Consultation with GENIE’s authors clarifies that this problem stems from the handling of binding energy during the NCQE process in GENIE. Note that the most recent GENIE versions are anticipated to improve the handling of binding energy in the modeling of the FSI processes. To address the issue, their recommendation involves marking a nucleon as intermediate if it lacks sufficient kinetic energy to leave a nucleus (e.g., its kinetic energy inside the nucleus is lower than the binding energy), with all its energy-momentum absorbed by the hadronic blob. Following these suggestions, the final-state information from the new GENIE models (i.e., Model-G for 2, 3, 4) is modified. To distinguish models subjected to this operation from those that are not, the modified models are re-labeled as Model-G for 2′, 3′, 4′.
@@
“New GENIE′ model variations”: the model variations among Model-G for 2′, 3′, 4′, which are used to demonstrate the impact of modifications in the new GENIE models, particularly addressing the issue of zero kinetic energy nucleons production in final-state results.
@@
When excluding nucleons with zero kinetic energies in the final-state, Model-G2′/4′ (N) exhibit a 25% lower rate of neutrons compared to the associated models. Similarly, Model-G3′ (A) shows a 10% lower rate of neutrons in comparison. Therefore, in the new GENIE′ models, Model-G2′ still exhibits the lowest rates, while Model-G3′ shows the highest rates.
@@
Fig. 5 presents the production rates of primary final-state particles, such as and , for kinetic energy. The dominant contributors to production are the QE and . To enhance clarity, nucleons with zero kinetic energy are excluded in Fig. 5. This operation results in kinetic energy spectra in Model-G 2, 3, 4 aligned with those from the corresponding Model-G 2′, 3′, 4′ in the left panels of Fig. 5, illustrating the spectra of and from NC interactions without deexcitation. When focusing on nucleon productions excluding those from deexcitation, we observe the same conclusions as depicted in Fig. 4. Furthermore, the impact of different nuclear models on production rates falls within the (0 20) MeV range in kinetic energy, as evident in the comparison between Model-G2 and Model-G4, or Model-N6 and Model-N7. Similarly, the effects of different FSI models on production rates are within the (0 40) MeV range in kinetic energy, observed in comparisons such as Model-G2 versus Model-G3, or Model-G2 versus Model-N6. By comparing the left and right panels of Fig. 5, it becomes apparent that the enhanced production rates of and from deexcitation occur within the (0 20) MeV range of kinetic energy. As expected, due to slight differences in the excited residual nucleus, the kinetic energy spectra of Model-G 2′, 3′, 4′ exhibit minor variations from the corresponding Model-G 2, 3, 4.
@@ 3.2.2 Neutron multiplicity
In Fig. 6, we examine the neutron multiplicity in the final-state of NC interactions involving -12C. Considering that the predominant neutron multiplicities are below three, the event rates for neutron multiplicities less than three are highlighted in the subpanels of the upper panels of Fig. 6. Due to the significant similarity in distributions between Model-G2 and Model-G4, as well as between Model-G2′ and Model-G4′, Model-G4/4′ is excluded from the illustrations. Excluding neutrons from the deexcitation process, the events with neutron multiplicities less than three account for approximately 91% in the new GENIE models using the N FSI model (i.e., Model-G2/4) and A FSI model (i.e., Model-G3). This calculation includes neutrons with zero kinetic energies in the final-state. If we exclude these neutrons in Model-G 2′, 3′, 4′, the events with neutron multiplicities less than three increase to 95% in Model-G2′/4′, while the fraction remains unchanged in Model-G3′. This difference arises because neutrons with zero kinetic energies from the N FSI model impact all neutron multiplicities, whereas those from the A FSI model only affect neutron multiplicities less than three. Compared to GENIE models, NuWro models produce around 99% events with neutron multiplicities less than three. For comparison purposes, the middle panels of Fig. 6 display the ratio of results related to multiplicities less than three to those in Model-G2. The neutrons induced by the deexcitation process increase the multiplicity. The parameter, , defined as the ratio of event rates for NC events with deexcitation to those for NC events without deexcitation, concerning multiplicities less than three is illustrated in the lowest panel of Fig. 6.
@@
Upon scrutinizing model variations, the maximum difference is observed in NC events associated with 10B for the new GENIE model variations. However, considering that nucleons with zero kinetic energy remain within the nuclei, the disparity in NC events with 10B is significantly mitigated in new GENIE′ model variations, resulting in lower event rates. Meanwhile, the event rate for NC events with 11C sees a noteworthy increase in new GENIE′ models. Regarding new NuWro model variations, the primary distinction arises in NC events with 11C, while other channels exhibit similarity in the absence of deexcitation scenarios. Notably, Model-N6 with the LFG nuclear model produces substantially higher event rates for NC events with 11C.
@@ 4 NC Background Predictions at Low Visible Energies
Fig. 10 presents event rates with different final-state nuclei across distinct prompt energy ranges of (0.7 100) MeV, (0.7 12) MeV, and (12 30) MeV, where the last two energy regions are associated with the detection regions of reactor neutrino and DSNB signals, respectively. The predominance of the NC background with 11C is evident. The deexcitation effect minimally impacts the overall event rate, causing a reduction in rates for dominant channels with 11C and 10B, while increasing rates for channels with lighter nuclei such as 9Be, 6Li and 8Be. When comparing the new and old neutrino interaction models, the rates of the new models do not surpass those of the old models, with only slightly higher rates observed for new GENIE′ models in the (0.7 12) MeV energy range.
Tab. 2 summarizes the average rates of the NC background rates, with the model uncertainty represented by the 1 deviation, for both old and new neutrino interaction models. In this table, the old models considered are Model-G1 and Model-N, where 1, 2, 3, 4, 5. Due to a strong correlation between the new GENIE and GENIE′ models, we opt to use the new GENIE′ model calculations instead of the new GENIE models for the prediction. Consequently, the results from the new models in Tab. 2 include Model-G, 2′, 3′, 4′, and Model-N, 6, 7. To investigate the impact of the deexcitation on the final NC background rate, both results without and with deexcitation are presented in Tab. 2. It’s important to note that while the deexcitation model used in this work still requires constraints from experimental data, the NC background rate with deexcitation is considered as the nominal result, as deexcitation always occurs in a realistic condition. In the prompt range of (0.7 100) MeV, the average NC background rates of the new models are lower than those of the old models for both scenarios, with and without the deexcitation process. This difference arises from the fact that the new models no longer incorporate the model with 1.35 GeV, which previously led to a higher NC background rate in the higher energy range. The deexcitation effect results in an increase in the average NC background rates by around a factor of 1.06 and 1.09 for old models and new models in this energy range, respectively. When considering the deexcitation, the associated uncertainty arising from model variation is around 16% and 14% for old and new models, respectively. In the (0.7 12) MeV and (12 30) MeV energy ranges, we will discuss the results related to the detection of reactor and DSNB in the LS detectors.
@@ Acknowledgements
The authors would like to thank Wan-Lei Guo and Xian-Guo Lu for helpful discussions. We also thank Davide Basilico and Liang Zhan for carefully reading the manuscript and useful comments. This work is supported in part by the National Key R&D Program of China under Grant No. 2024YFE0110500, by National Natural Science Foundation of China under Grand Nos. 12405125, 12075255, 11835013, 12125506.